OpenStudy (anonymous):

another integral \[\int \frac{1}{y^{2}+y} dy\]

OpenStudy (anonymous):

sorry I feel like I should know this

OpenStudy (anonymous):

partial fractions.... find A and B in the decomposition for the integrand: \[\huge \frac{1}{y^2+y}=\frac{A}{y}+\frac{B}{y+1} \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did you do this before?

OpenStudy (anonymous):

I believe so, lets see if I remember

OpenStudy (anonymous):

you'll end up with: \[\huge A(y+1)+By=1 \]

OpenStudy (anonymous):

\[\huge \frac{A(y+1)y}{y}+\frac{B(y+1)y}{y+1}= 1\]

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

now all you gotta do is to choose "nice" values of y to make it easy in finding values for A and B... for example, y=-1... solve for B... B=???

OpenStudy (anonymous):

\[\huge A(y+1)+By=1\] -1

OpenStudy (anonymous):

B=-1 you mean right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok... so what other "nice" value for y will make it easy to find A?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

yes... so A=??

OpenStudy (anonymous):

1

OpenStudy (anonymous):

there is also another way to do it without guessing for y values right?

OpenStudy (anonymous):

good... so now you have: \[\huge \frac{1}{y^2+y}=\frac{1}{y}-\frac{1}{y+1} \]

OpenStudy (anonymous):

yes, thanks!

OpenStudy (anonymous):

hmm... they weren't actually guesses... they were "convenient" values for y...

OpenStudy (anonymous):

I know, but I thought there was a way. A+B=... Oh wait I think I remember... one sec

OpenStudy (anonymous):

oh... i think i know what you're getting at.....

OpenStudy (anonymous):

You can either do test cases or solve a system of equations or some of both methods. :-)

OpenStudy (anonymous):

and yes.... if that denominator was something more complicated, you'd have to solve a system...

OpenStudy (anonymous):

@dpaInc was doing test cases as much as I can tell from skimming

OpenStudy (anonymous):

\[\huge A(y+1)+By=1\] how would I do that again?

OpenStudy (anonymous):

"convenient" values for y.. if y=-1, A will dissapear... if y=0, B will disappear...

OpenStudy (anonymous):

I guess the problem isn't complicated enough to solve a system

OpenStudy (anonymous):

I understand how we solved this problem, by eliminating A to solve for B and vice versa

OpenStudy (anonymous):

no.... but even when doing the system, if i remember correctly you'd still have to do this "convenient" thingy...

OpenStudy (anonymous):

makes sense

OpenStudy (anonymous):

thanks again @dpaInc !

OpenStudy (anonymous):

yw... seems like you've gone through all this before... are you reviewing?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

nice work....:)

OpenStudy (anonymous):

I guess I have forgotten almost everything about solving integrals

OpenStudy (anonymous):

they'll come back to you....:)

OpenStudy (anonymous):

I'm patiently waiting for that to happen LOL! thanks :)

OpenStudy (anonymous):

:)